# Training-free Camera Control for Video Generation

### v3

@@ Appendix C Visual Comparisons with State-of-the-arts
Sec. 3 showcases some qualitative comparisons between CamTrol and different state-of-the-art approaches. For better visualization and comparison, we provide more video results in the supplementary materials. Please check them if needed.
@@ Appendix D Implementation Details
For text prompt input, we use Stable Diffusion v2-1  [^2] https://huggingface.co/stabilityai/stable-diffusion-2-1 or Stable Diffusion XL  [^3] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0 to generate the initial image. The inpainting model we apply is Stable Diffusion inpainting model proposed by runway  [^4] https://huggingface.co/runwayml/stable-diffusion-inpainting, and backward step of inpainting is set to 25. We deploy ZeoDepth  [^5] https://github.com/isl-org/ZoeDepth as depth estimation model. The classifier-free guidance scale follows default setting of base models themselves. We use SVD’s default setting of 6 fps for video generation, and process reference videos to 6 fps for FVD and FID calculation. The complete procedure of CamTrol is elucidated in Algorithm 1.
For computational analysis, we set both number of frames and decoding size to 14, generation steps to 25. We do not apply xformers in all approaches. The computational analysis at $576\times 1024$ resolution is shown in Table 4. SVD-based CameraCtrl only support resolution at $320\times 576$.
@@
SVD MotionCtrl CameraCtrl CamTrol ($t_{0}=10$)
@@
Table 5: Definitions of camera motions. We follow the terminology in cinematography to describe different camera movements.
@@
Pedestal Up Moving a camera vertically in its entirety.
@@
Truck Left Moving a camera horizontally in its entirety.
@@
Roll Clockwise Rotating a camera in its entirety in a horizontal manner.
@@
Rotate Clockwise Moving a camera around a subject.
@@
Hybrid Combinations Spatial and temporal combination of basic motions.
Complicated Arbitrary Camera extrinsic matrix or a reference video as input.
@@ 12 Sample random noise $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ ;
In Sec. 3.1, we designate point cloud as the intermediate 3D representation for explicit camera motion modeling. Doubts may arise why we do not use more complex 3D representation which might be more useful. Here we take the most prevalent 3D representation: 3D Gaussian Splatting  [^6] https://repo-sam.inria.fr/fungraph/3d-gaussian-splatting/, as example to elaborate the benefit of using point cloud in three aspects:
Firstly, concerning the input, point cloud is able to construct the whole scene from one single input image combining techniques of depth estimation and inpainting. GS, though also an explicit 3D representation, requires optimization following images from different views, which means it is neither capable of handling single input image, nor can it leverages 2D inpainting to complete the scene gradually.
Secondly, from the aspect of time, point cloud can directly lift 2D points onto 3D spaces, while 3DGS demands optimization on each scenario. As a training-free method, our method takes nearly no time to generate a video after multi-view images are acquired, but would need more time if 3DGS were applied.
Lastly, from the task itself, the target of stage I is not precisely reconstruct the 3D scene but only offer a rough layout guidance, in this context, rendered images from point cloud qualify enough and no further refinement on 3D reconstruction is necessary.
From the analysis, point cloud is quite qualified serving as a rough layout guidance in a relatively quick speed, without any further optimization and redundant multi-view images as input. The use of point cloud allows our algorithm to be totally training-free and optimization-free, simultaneously being able to produce camera-moving videos quickly with merely one single image or text prompt as input.
@@ Appendix F Details about Generating 3D Rotation Videos
3D generation models (Voleti et al., 2024; Melas-Kyriazi et al., 2024; Shi et al., 2023; Chen et al., 2024; Han et al., 2024) are trained on highly-regulated 3D datasets, these datasets are hard to collect, and consist only a limited variety of data types (e.g. single static objects with no background). As a consequence, 3D generation models exhibit very constrained output distribution and could only produce results in certain styles. CamTrol avoid these problems of arduous data collecting, laborious finetuning and output collapse by utilizing the layout prior hidden in video foundation models. Not only does it require no training, but this advantage also benefits CamTrol from inheriting most of the prior knowledge inside video foundation models, e.g., the diversity of scenarios, the dynamic of moving objects, temporal consistency, etc. Thus CamTrol is able to generate dynamic 3D content, both objects and scenes, in a totally unsupervised and training-free manner, this is what other method cannot achieve yet.
Compared with regulated datasets, the problem of processing wild pictures in 3D is that some of the parameters is unknown. In Sec. 3.1, we’ve mentioned that the camera intrinsic matrix $\mathbf{K}$ and initial extrinsic matrix $\mathbf{P}_{0}$ are set by convention as they’re usually intractable. Another crucial parameter concerning 3D video generation is the distance between camera and the content of input image (denote as $f$), note that input image could be synthetic or real. Considering most camera rotations are done around the center point, we extract a patch from the very center of the input image, conducting depth estimations on the it and define the distance $f$ as the averaged depth. The rotation and transition matrix during camera moving can be formed as :
@@
Here $i\in[0,N-1]$ and $\theta_{i}$ refers to rotation angle around $y$ axis at step $i$, $D$ denotes depth estimation of image, $P$ represents the patch around central point $(x_{0},y_{0})$. In our experiment, we choose $(j,k)\in[-10,10]$ as the size of patch.
@@ Appendix G Motion Blur, Problems and Solutions
Videos produced by CamTrol need to satisfy certain camera movement, and some drastic perspective changes might cause visible trails, recognised as motion blur of objects or scenes. This phenomenon will appear to be more indispensable when video base model holds a relatively small generation length (e.g. 16 frames) as well as the motion scale becomes larger (e.g. tilt up for 60 degrees or more). To avoid blur issues in controlling video camera motion, we propose several solutions as below:
@@
    According to the analysis above, the blur issue is caused by limited generation frames and large camera movement, thus the most intuitive solution is to either cut down motion scale or utilize a more capable generation model. For severe perspective changes, the optimal approach would involve employ video foundation models that support larger generation length (e.g. CogVideoX (Yang et al., 2024b) supports generating video with 49 frames). This allows model to manage motions of equivalent magnitude while experiencing a smaller moving range between adjacent frames, thereby brings effective alleviation to blur problems.
@@
    One can also stack the results of multiple generations to form a complete outcome, i.e., treating the last frame of previous generation as the starting frame for the next, and integrate them as a whole. This approach is more suitable when utilizing a image-to-video (I2V) base model. Since most open-source video foundation models are text-to-video (T2V), one may consider increasing the step of camera motion inversion $t_{0}$, which guarantees more fidelity towards input images’ content (and motion).
@@
    Besides the above two approaches, it is as well a common and convenient choice to apply frame interpolation towards the output. Lots of off-the-shelf frame interpolation models are open-sourced and could be found on github.
@@
    Lastly, if video length is not being able to alter, it may be necessary to increase the fps of generated videos for better visualization. Although larger fps leads to shorter video duration, it simultaneously makes the visual persistence brought by motions less pronounced, which reduce blur visually.
In our experiments, we take raw output in all settings.

### v4

@@ Appendix C Visual Comparisons with State-of-the-arts
Sec. 3 showcases some qualitative comparisons between CamTrol and different state-of-the-art approaches. For better visualization and comparison, we provide more video results in the supplementary materials.
@@ Appendix D Implementation Details
For text prompt input, we use Stable Diffusion v2-1  [^2] https://huggingface.co/stabilityai/stable-diffusion-2-1 or Stable Diffusion XL  [^3] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0 to generate the initial image. The inpainting model we apply is Stable Diffusion inpainting model proposed by Runway  [^4] https://huggingface.co/runwayml/stable-diffusion-inpainting, and the backward step of inpainting is set to 25. We use ZeoDepth  [^5] https://github.com/isl-org/ZoeDepth as depth estimation model. The classifier-free guidance scale follows the default setting of the base models themselves. We use SVD’s default setting of 6 fps for video generation, and process reference videos to 6 fps for FVD and FID calculations. The complete procedure of CamTrol is elucidated in Algorithm 1.
For computational analysis, we set both the number of frames and the decoding size to 14, and the generation steps to 25. We do not apply xformers in any of the approaches. The computational analysis at $576\times 1024$ resolution is shown in Table 4. SVD-based CameraCtrl only supports a resolution at $320\times 576$.
@@
SVD MotionCtrl CameraCtrl CamTrol ($t_{0}=15$)
@@
Table 5: Definitions of camera motions. We adopt the terminology from cinematography to describe different camera movements.
@@
Pedestal Up Moving the camera vertically in its entirety.
@@
Truck Left Moving the camera horizontally in its entirety.
@@
Roll Clockwise Rotating the camera in its entirety in a horizontal manner.
@@
Rotate Clockwise Moving the camera around a subject.
@@
Hybrid Combinations Spatial and temporal combinations of basic motions.
Complicated Arbitrary Camera extrinsic matrices or reference videos as input.
@@ 12 Sample random noise $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ ;
In Sec. 3.1, we designate point cloud as the intermediate 3D representation for explicit camera motion modeling. Doubts may arise as to why we do not use a more complex 3D representation which might be more useful. Here we take the most prevalent 3D representation–3D Gaussian Splatting  [^6] https://repo-sam.inria.fr/fungraph/3d-gaussian-splatting/, as an example to elaborate on the benefits of using point cloud in three aspects:
Firstly, concerning the input, point cloud is able to construct the entire scene from one single input image combining techniques of depth estimation and inpainting. GS, though also an explicit 3D representation, requires optimization based on images from different views, which means it is neither capable of handling single input image, nor can it leverage 2D inpainting to gradually complete the scene.
Secondly, from the aspect of time, point cloud can directly lift 2D points onto 3D spaces, while 3DGS demands optimization for each scenario. As a training-free approach, our method takes almost no time to generate a video after multi-view images are acquired, but would need more time if 3DGS were applied.
Lastly, from the task itself, the goal of stage I is not to precisely reconstruct the 3D scene but rather to offer a rough layout guidance. In this context, rendered images from point cloud are sufficiently qualified, and no further refinement of the 3D reconstruction is necessary.
From the analysis, point cloud is quite qualified serving as a rough layout guidance in a relatively quick speed, without any further optimization or redundant multi-view images as input. The use of point cloud allows our algorithm to be completely training-free, while still being able to produce camera-moving videos quickly with only one single image or text prompt as input.
@@ Appendix F Details about Generating 3D Rotation Videos
3D generation models (Voleti et al., 2024; Melas-Kyriazi et al., 2024; Shi et al., 2023; Chen et al., 2024; Han et al., 2024) are trained on highly-regulated 3D datasets, these datasets are hard to collect, and include only a limited variety of data types (e.g., single static objects with no background). As a consequence, 3D generation models exhibit a very constrained output distribution and can only produce results in certain styles. CamTrol avoids these problems of arduous data collection, laborious finetuning, and output collapse by utilizing the layout prior hidden in video foundation models. Not only does it require no training, but this advantage also benefits CamTrol from inheriting most of the prior knowledge inside video foundation models, such as the diversity of scenarios, the dynamics of moving objects, temporal consistency, etc. Thus, CamTrol is able to generate dynamic 3D content, both objects and scenes, in a totally unsupervised and training-free manner, this is what other methods cannot achieve yet.
Compared with regulated datasets, the problem of processing wild pictures in 3D is that some of the parameters are unknown. In Sec. 3.1, we mentioned that the camera intrinsic matrix $\mathbf{K}$ and the initial extrinsic matrix $\mathbf{P}_{0}$ are set by convention, as they are usually intractable. Another crucial parameter concerning 3D video generation is the distance between the camera and the content of input image (denoted as $f$), note that input image could be synthetic or real. Considering that most camera rotations are performed around the center point, we extract a patch from the very center of the input image, conduct depth estimations on it, and define the distance $f$ as the average depth. The rotation and translation matrices of camera movement can be formed as:
@@
Here, $i\in[0,N-1]$ and $\theta_{i}$ refer to the rotation angle around the $y$ axis at step $i$, $D$ denotes the depth estimation of the image, and $P$ represents the patch around the central point $(x_{0},y_{0})$. In our experiment, we choose $(j,k)\in[-10,10]$ as the size of the patch.
@@ Appendix G Motion Blur, Problems and Solutions
Videos produced by CamTrol need to satisfy certain camera movements, and some drastic perspective changes might cause visible trails, recognized as motion blur of objects or scenes. This phenomenon will appear to be more indispensable when the video base model has a relatively small generation length (e.g., 16 frames) as well as the motion scale becomes larger (e.g., tilt up by 60 degrees or more). To avoid blur issues when controlling video camera motion, we propose several solutions as follows:
@@
    According to the analysis above, the blur issue is caused by limited generation frames and large camera movements, thus the most intuitive solution is to either reduce motion scale or use a more capable generation model. For severe perspective changes, the optimal approach would involve employing video foundation models that support larger generation lengths (e.g., CogVideoX (Yang et al., 2024b) supports generating videos with 49 frames). This allows the model to manage motions of equivalent magnitude while experiencing a smaller moving range between adjacent frames, thereby bringing effective alleviation to the blur problem.
@@
    One can also stack the results of multiple generations to form a complete output, i.e., treating the last frame of the previous generation as the starting frame for the next and integrating them as a whole. This approach is more suitable when using an image-to-video (I2V) base model. Since most open-source video foundation models are text-to-video (T2V), one may consider increasing the step of camera motion inversion $t_{0}$, which guarantees more fidelity towards input images’ content (and motion).
@@
    Besides the above two approaches, applying frame interpolation to the output is also a common and convenient choice. Many off-the-shelf frame interpolation models are open-source and can be found on GitHub.
@@
    Lastly, if the video length cannot be altered, it may be necessary to increase the fps of the generated videos for better visual quality. Although a larger frame rate leads to a shorter video duration, it simultaneously makes the visual persistence brought by motions less pronounced, which reduces blur visually.
In our experiments, we take the raw output in all settings.